### v3

# Adaptive NAD: Online and Self-adaptive Unsupervised Network Anomaly Detector

Abstract
The widespread usage of the Internet of Things (IoT) has raised the risks of cyber threats, thus developing Anomaly Detection Systems (ADSs) that can adapt to evolving or new attacks is critical. Previous studies primarily focused on offline unsupervised learning methods to safeguard ADSs, which is not applicable in practical real-world applications. Besides, most of them strongly rely on assumptions of known legitimates and fail to satisfy the interpretable requirements in security applications, creating barriers to the adoption in practice. In this paper, we design Adaptive NAD, a general framework to improve and interpret online and Adaptive unsupervised Network Anomaly Detection in security domains. An interpretable two-layer anomaly detection strategy is proposed to generate reliable high-confidence pseudo-labels. Then, an online learning scheme is introduced to update Adaptive NAD by a novel threshold calculation technique to adapt to new threats. Experimental results demonstrate that Adaptive NAD achieves more than 5.4%, 23.0%, and 3.2% improvements in SPAUC compared with state-of-the-art solutions on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively. The code is released at https://github.com/MyLearnCodeSpace/Adaptive-NAD.
Index Terms:
Anomaly detection, unsupervised learning, interpretability, online learning, dynamic threshold
I Introduction
The rapid expansion of the Internet of Things (IoT) has revolutionized many areas, such as healthcare, manufacturing, and autonomous driving [1, 2, 3]. However, it also increases risks of cybersecurity attacks, for example, UnitedHealth’s \$872 Million Cyberattack in 2024  [^1] https://www.sec.gov/Archives/edgar/data/731766/000073176624000146/a2024q1exhibit991.htm and the I.M.F. breach in 2024  [^2] https://www.bleepingcomputer.com/news/security/international-monetary-fund-email-accounts-hacked-in-cyberattack/. Anomaly Detection Systems (ADSs) are essential to discover potential network security problems. The advances in machine learning, particularly deep learning, have significantly accelerated this progress [4, 5, 6, 7].
In the real world, the environment is dynamic and ever-changing. Both normal data and anomalies are evolving over time. Thus, traditional offline training-based methods are not applicable in practical applications. Given that machine learning models strongly rely on learned data patterns, any shifts in normal/abnormal behavior could damage anomaly detection performance. Hence, ADSs demand constant adjustment to the model’s changing behavior over time in an online manner. Online learning provides a potential way to improve models’ adaptability by continuously retraining the model. Moreover, unsupervised learning-based models can directly learn from only normal data without any known threats, they are more favorable for online learning tasks of anomaly detection due to the inherent lack of labeled anomalies in historical data and the unpredictable and highly varied nature of anomalies.
In practical security-related applications, it is a challenge to establish an effective online unsupervised learning system demanding expeditious adaptation of the model to a continuous incoming of unlabeled data.
Primarily, existing unsupervised anomaly detection approaches strongly rely on assumptions of known legitimate [8, 9, 10, 11, 12, 13, 14, 11, 15, 16, 17] and/or even malicious data [18, 19] to support the training of the initial unsupervised model. Moreover, recent existing unsupervised-learning-based research [8, 9, 10, 20, 15] conducts anomaly detection by using a fixed threshold, which tends to fail to detect anomalies when concept drifts exist. To address this, a dynamic threshold is utilized to distinguish normal from anomalies in [11, 21], but suffers from high false positive rates. Additionally, most of existing literature [8, 9, 10, 12, 14, 15, 22, 16] trains their unsupervised models offline without any adaption to drifting patterns, making them not applicable in real-world applications. There is only limited research on online unsupervised anomaly detection, and most of them like [23, 24, 25] consider “online” as the model being trained offline and utilized for real-time inferences without model updating.
Furthermore, since either online or offline unsupervised deep learning models, like LSTM, are not transparent and their prediction results are not interpretable, it is hard to trust the system’s decision (e.g., normal events or anomalies) without sufficient reasons and credible evidence in security-related applications, especially in healthcare, manufacturing, or autonomous driving [1, 2, 3]. However, most of the existing unsupervised anomaly detection models like [8, 9, 10, 11, 12, 13, 11, 15, 16, 17] lack interpretability. Although some work [25, 14] provides some interpretability by comparing the reconstruction error or through the learned graph edges and attention weights, there are still some research gaps in developing interpretable online unsupervised anomaly detection models in ADSs.
Figure 1: The proposed Adaptive NAD framework.
Our work. To address the above issues, we propose a general framework, named Adaptive NAD, to interpret and improve online unsupervised deep learning-based anomaly detection in security applications. The high-level goal of Adaptive NAD is to design a novel interpretable online unsupervised framework for anomaly detection that meets several practical security requirements (such as generalization, human-readable, low false positives, and adapting new threats). To this end, we designed two key strategies in Adaptive NAD: (i) interpretable two-layer based unsupervised anomaly detection strategy by leveraging advantages of both supervised and unsupervised models to generate highly confident pseudo labels in highly imbalanced network streaming data. (ii) Online learning strategy for fine-tuning models and regenerating loss distributions aligned with the evolving network events to alleviate the effect of concept drift.
Overall, Adaptive NAD is a novel online and self-adaptive unsupervised network anomaly detection framework with interpretable ability, that can be utilized with any pair of unsupervised-supervised models for online unsupervised anomaly detection tasks without any knowledge of both normal and abnormal events.
The contributions of this paper are mainly three-fold:
  • $\bullet$
    A novel interpretable two-layer structure that can be utilized with any pair of unsupervised and supervised learning models for online unsupervised anomaly detection without any label information of normal events and anomalies.
  • $\bullet$
    A new dynamic threshold technique with theoretical proof for determining high-confidence pseudo labels can be employed for fine-tuning online models.
  • $\bullet$
    We compare it with eleven state-of-the-art models for example, GDN [14], DAGMM [22], AOC-IDS [26], DeepFT [11], DeepAID [15], and Online-Offline Framework (OOF) [21] on two classical public anomaly detection datasets (CIC-Darknet2020 [27], CIC-DoHBrw-2020 [28], and Edge-IIoTset [29]) and the results show more than 5.4%, 23.0%, and 3.2% 3335.4%, 23.0%, and 3.2% are the improvements compared with the best state-of-the-art. improvements in SPAUC.
The rest of the paper is outlined as follows: we first give an overview of the system design in Section II, including design goals, system components, and details of the models used in Adaptive NAD. Then, the proposed novel techniques, i.e., interpretable two-layer anomaly detection strategy (ITAD-S), online learning scheme, and threshold calculation technique, are presented in Section III along with a comprehensive complexity and storage analysis of the proposed Adaptive NAD framework. Following it, Section IV describes the experimental setup and results. Existing related literature is discussed in Section VI. Finally, we conclude the paper in Section VII.
II System Design
This section presents the design goals, system components of the proposed Adaptive NAD, and details of the two key models utilized in Adaptive NAD.
II-A Design Goals
The design goals of our proposed approach are explained below:
  1. 1.
    Capable of handling diverse data: System networks collected from different devices vary greatly regarding their format and content. A generalized ADS needs to be capable of analyzing various network formats.
  2. 2.
    Capable of adapting to new cyber threats: System behavior and potential threats evolve, and consequently, so do the system networks and attack models. The ability of ADS to learn newer event patterns should be automated without requiring the experts’ effort.
  3. 3.
    High accuracy but low false positive rate: Anomaly detection should correctly predict both normal and abnormal activities; in other words, achieving a high accuracy rate but a low false positive rate.
  4. 4.
    Low computational cost: High computational cost incurred during anomaly detection leads to increased latency, deployment, and maintenance costs.
II-B System Components
As shown in Fig. 1, the proposed system comprises the following components:
  • $\bullet$
    User Devices: This module stores and feeds network data to the unsupervised deep learning model $M$ in the server. The networks are collected from multiple distributed users’ devices in the network and contain information about different system states and events of interest defined by the system administrators.
  • $\bullet$
    Unsupervised deep Learning Model $M$: This model is used to learn the network traffic using unsupervised deep learning models. In this paper, an unsupervised learning model (i.e., NAD) is utilized to learn the pattern of the normal network traffic (see Section II-C).
  • $\bullet$
    Predictor: After the deep learning model $M$ is trained on the normal network traffic, $M$ can be used to predict the new upcoming network data.
  • $\bullet$
    Adaptive Threshold Computation: This model generates two thresholds $T_{1}$ and $T_{2}$ by using Quantiles and Maximum Likelihood Estimation (MLE) (see Section III-C). Thresholds are adaptively changing to incorporate the characteristics of new network traffic to identify normal and abnormal network data.
  • $\bullet$
    Decision: This module is employed to select network data with the (1-$\alpha$) confidence level by dynamic thresholds, for the training of the machine learning model $M^{{}^{\prime}}$ based on the predicted network loss data from the unsupervised deep learning model $M$. In addition, the loss of accurately selected network events is utilized to update the thresholds.
  • $\bullet$
    Supervised Machine Learning Model: This module uses simple machine learning methods, such as Random Forest classifier (RF), to predict the uncertainty network traffic from an unsupervised deep learning model. This improves the detection performance of unsupervised learning models significantly.
  • $\bullet$
    Database: The stored dataset is used to update the machine learning model.
II-C The Network Anomaly Detection Model (NAD)
The Network Anomaly Detection Model (NAD) is utilized as the unsupervised deep learning model $M$, which is adapted from VAE [30]. We use the Long Short-Term Memory (LSTM) for the encoder and decoder phase and use a Gaussian distribution for $p(z)$ in VAE. NAD is a generative model, whose aim is to perform and learn a probabilistic model $P$ from a sample $X$ with unknown distribution. It overcomes the three following drawbacks from its predecessors: Firstly, it makes strong assumptions about the structure of the data. Secondly, due to the approximation methods, they end up with sub-optimal models. Lastly, they generally use significant computational resources for the inference procedure like Markov Chain Monte Carlo (MCMC) [31].
The idea behind this model is to use latent variables $Z$, which might be understood as hidden variables that influence the behavior of $X$; thus, having the initial sample $X$, we need to train a model to learn the distribution of $Z$ given the distribution $X$. This allows us to eventually make the inference over $X$. We now express the model in probability terms, as follows. Specifically, for $x\in X$, $z\in Z$:
[…]
Since the integral with respect to $z$ in the Eq. (1) is difficult to compute due to the high dimensional space of $Z$, we can express the log maximum likelihood of $logp_{\theta}(x)$ as [30]:
$logp_{\theta}(x)=D_{kl}[p_{\theta}(z|x)||p_{\theta}(z)]+\mathcal{L}(x,\phi,\theta),$ (2)
where in this equation, the first term $D_{kl}$ is the Kullback-Leibler divergence (a non negative term) and the second term can be seen as a lower bound of likelihood $logp_{\theta}(x)$. Thus, this bound is referred to as the variational lower bound, which can be expressed as: [30]:
$logp_{\theta}(x)\geq\mathcal{L}(x,\phi,\theta)=E[logp_{\theta}(x,z)-logp(z|x)].$ (3)
Due to the intractability of $p(z|x)$, NAD uses a learning model $q_{\phi}(z|x)$ that approximates $p(z|x)$ with $\phi$ as the parameter of distribution $q$. Then, we can rewrite the equation as:
$\mathcal{L}(x,\phi,\theta)=E_{q_{\phi}(z|x)}[logp_{\theta}(x,z)-logq_{\phi}(z|x)]$
$=-D\_kl[q_{\phi}(z|x)||p_{\theta}(z)]+E_{q_{\phi}(z|x)}[logp_{\theta}(x|z)],$ (4)
where $p(z)$ is in general a simple distribution like Gaussian for continuous variable or Bernoulli for binary variable. Finally, we want to maximize $logp_{\theta}(x)$, but in Eq. (4) we observe that we need to minimize the Kullback-Leibler divergence. Thus, NAD optimizes the function $\mathcal{L}$ for this purpose [30].
Fig. 2(a) shows the NAD diagram. In its Encoder phase, we can see that given an input $X$, the model uses the LSTM encoder $q_{\phi}(z|x)$ with the multilayer network $f(x,\phi)$ to approximate the probability $p(z/x)$, which is the distribution of the latent variable $Z$ given $X$. In its Decoder phase, with the knowledge of $Z$, the model uses a LSTM decoder $p_{\theta}(x|z)$ with the multilayer network $g(z,\theta)$ to approximate the probability $p(x/z)$, which is the distribution of the input variable $X$ given its latent variable $Z$. Finally, the NAD model solves the optimization problem in Eq. (3) using the Encoder and Decoder learning models given as Eq. (5) and Eq. (6), respectively.
$q_{\phi}(z|x)=q(z,f(x,\phi)),$ (5)
$p_{\theta}(x|z)=p(x,g(x,\theta)).$ (6)
(a) NAD Event Model.
(b) RF classifier.
Figure 2: Selected models.
II-D Random Forest
Here, we use the Random Forest algorithm [32] as our supervised machine learning model $M^{\prime}$, which assumes that $X_{1},X_{2},\cdots,X_{M}$ are $M$ attributes. A sample $S$ is denoted as a vector $(x_{1},x_{2},\cdots,x_{M})$, where $x_{i}$ is the observed value of $X_{i}$. We also define the class label $C=\{c_{1},c_{2}\}$ consisting of two different values $c_{1}$ (normal) and $c_{2}$ (abnormal) whose classification depends on the attributes. Hence, the class predicted by the Random Forest method is given as,
$C_{rf}(S)=\operatorname*{arg\,max}_{c\in C}p(c)\prod_{i}p(x_{i}|c),$
where training samples are employed to predict the $p(x_{i}|c)$ value. The decision boundary, the curve that separates the points in two groups $c_{1}$ and $c_{2}$, is calculated using the classifier function $f_{rf}$:
$f_{rf}(S)=\frac{p(c=c_{1})}{p(c=c_{2})}\prod_{i}\frac{p(x_{i}|c=c_{1})}{p(x_{i}|c=c_{2})}$ (7)
Hence, the network traffic belongs to class $c_{1}$ if $f_{rf}(S)\geq 1$, $C_{2}$ otherwise. Fig. 2(b) shows a conceptual view of the RF classifier.
III Methodologies
III-A Interpretable Two-layer Anomaly Detection Strategy (ITAD-S)
ITAD-S consists of a two-layer model, i.e., an Unsupervised Deep Learning model $M$ in the first layer and an interpretable Supervised Machine Learning model $M^{\prime}$ [^4] In this paper, we use LSTM-VAE and RF as an example of the unsupervised deep learning model $M$ and interpretable supervised machine learning model $M^{\prime}$ to better elaborate our work. in the second layer (as shown in Fig. 1). The key idea behind it is that it skillfully leverages the advantages of supervised learning to improve the performance of unsupervised learning models and ultimately construct an efficient interpretable unsupervised learning framework for online anomaly detection in dynamic environments. This strategy has two critical processes, i.e., a high-confidence pseudo label generation process and a prediction process for fully automated unsupervised anomaly detection.
High-confidence pseudo label generation process: This process utilizes a novel Adaptive Threshold Calculation technique to select high-confidence labeled samples predicted by the unsupervised learning model from the input data. It is utilized to eliminate manual labeling for autonomous updates. Our framework does not assume any known normal or anomalous data. Given the fact that anomalies are usually much less than normal samples in practical security-related applications.
Specifically, we can use a limited dataset (e.g., the first day’s data or randomly select 1% from the dataset) as our first-round training set to train the unsupervised model $M$. As aforementioned, the upcoming network events are predicted by the trained unsupervised model $M$ to predict their loss values. The network events are then divided into three categories by predetermined thresholds $\{T_{1},T_{2}\}$ [^5] T1 and T2 are determined by the quantiles of the loss distributions for normal and abnormal samples, respectively. (where $T_{1}<T_{2}$): high-confidence anomalous samples (i.e., true positive network events), high-confidence normal samples (i.e., true negative network events), and uncertain samples, as illustrated in Fig. 1.
Prediction process.: This process predicts anomalies by the interpretable machine learning $M^{\prime}$ trained with the pseudo-labeled high-confidence samples from the previous process. The combination of both the high-confidence pseudo label generation process and the prediction process enhances the adaptive NAD framework’s learning ability and ensures a low false positive of the models. It is worth noting that the interpretable machine learning model $M^{\prime}$ (i.e., RF in our work) is employed to make final predictions, and once it detects an abnormal network event, it raises an alert immediately and sends this abnormal data to the system management.
Figure 3: Timeline of Adaptive NAD.
III-B Online Learning Scheme
The online learning scheme is introduced to periodically update thresholds $\{T_{1},T_{2}\}$ and the models $\{M,M^{\prime}\}$ in Adaptive NAD, enabling it to efficiently and accurately predict emerging anomalies (e.g., zero-day attacks) and evolving normal samples (e.g., new normal patterns) in complex real-world environments.
Input : Dataset for the first-round of training: $X_{0}$; Streaming new input vector: $\{x_{l_{0}+1},\ldots,x_{i},\ldots\}$, $i>l_{0},i\in\mathbb{N}$.
Output : Well-trained model $M$, $M^{\prime}$.
1 Randomly initialize the unsupervised model $M$
2 Randomly initialize the machine learning model $M^{\prime}$;
3 Train($X_{0},M$);
4 foreach $x_{0i}\in X_{0}$ do
5       $\text{Loss}\leftarrow M(x_{0i})$;
6       $LossDB_{\text{normal}}\leftarrow LossDB_{\text{normal}}\cup\{\text{Loss}_{0i}\}$;
[…]
8while inputting $x_{i}$ do
9       $\text{Loss}\leftarrow M(x_{i})$;
10       $BatchDB\leftarrow BatchDB\cup\{x_{i},\text{Loss}\}$;
11      if $\text{Size}(LossDB_{\text{abnormal}})<n$ then
12             $X\leftarrow X\cup\{x_{i}\}$;
13             if $\text{Loss}<T_{1}$ then
14                   $DB_{\text{normal}}\leftarrow LossDB_{\text{normal}}\cup\{\text{Loss}\}$ Loss };
15                   $Y\leftarrow Y\cup\{\text{normal}\}$;
16                   $X_{\text{normal}}\leftarrow X_{\text{normal}}\cup\{x_{i}\}$;
17            else
18                   $DB_{\text{abnormal}}\leftarrow LossDB_{\text{abnormal}}\cup\{\text{Loss}\}$ Loss };
19                   $Y\leftarrow Y\cup\{\text{abnormal}\}$;
20            continue;
21
22      if $\text{Loss}<T_{1}$ then
23             $DB_{\text{normal}}\leftarrow LossDB_{\text{normal}}\cup\{\text{Loss}\}$ Loss };
24             $X\leftarrow X\cup\{x_{i}\}$;
25             $Y\leftarrow Y\cup\{\text{normal}\}$;
26             $X_{\text{normal}}\leftarrow X_{\text{normal}}\cup\{x_{i}\}$;
27      else if $\text{Loss}>T_{2}$ then
28             $DB_{\text{abnormal}}\leftarrow LossDB_{\text{abnormal}}\cup\{\text{Loss}\}$ Loss };
29             $X\leftarrow X\cup\{x_{i}\}$;
30             $Y\leftarrow Y\cup\{\text{abnormal}\}$;
31      else if $T_{1}\leq\text{Loss}\leq T_{2}$ then
32             $y_{i}\leftarrow M^{\prime}(x_{i})$;
33             if $y_{i}==normal$ then
34                   $X_{\text{normal}}\leftarrow X_{\text{normal}}\cup\{x_{i}\}$;
35
36      if $\text{Size}(BatchDB)==m$ then
37             $T_{1}\leftarrow\text{AdaptiveThreshold}(LossDB_{\text{normal}},p_{\text{0}})$;
38             $T_{2}\leftarrow\text{AdaptiveThreshold}(LossDB_{\text{abnormal}},p_{\text{1}})$;
39             Train($X_{\text{normal}},M$);
40             Train($X,Y,M^{\prime}$);
41             $Empty(X,Y,X_{\text{normal}},BatchDB)$.
42
Algorithm 1 Online Learning Framework
Input : The database storing the loss of samples: $LossDB$.
Percentile for distribution: $p$.
Output : Threshold T.
1 if $\text{Size}(LossDB_{\text{normal}})>0$ then
2       $FitDist\leftarrow\text{FitBestDistribution}(LossDB)$;
[…]
4
5else
6       $\text{T}\leftarrow\emptyset$.
Algorithm 2 Adaptive Threshold
As depicted in Fig. 3, the proposed online learning framework follows a timeline that incrementally enlarges the training set with high-confidence pseudo-labels and regularly updates the unsupervised deep learning model $M$ with the high-confidence pseudo-labeled dataset. For example, given time series samples from $t=i$ to $t=i+2$, the loss values are first predicted/inferred using an unsupervised model. Based on these loss values and the predefined thresholds $\{T_{1},T_{2}\}$, high-confidence pseudo-labels are assigned to the samples. The samples and their corresponding labels are then used to update both the supervised and unsupervised models. Simultaneously, the thresholds $\{T_{1},T_{2}\}$ are updated with the new loss values. The updated models and thresholds will be used for predicting the samples in the next time period.
The online learning framework is presented in Algorithm 1, where $X$ and $Y$ represent high-confidence samples classified by the unsupervised model, $X_{\text{normal}}$ represents the normal samples used to train the unsupervised model, $n$ represents the number of training samples in the initial phase, $LossDB$ is a fixed-size, first-in-first-out (FIFO) queue, and $m$ represents the number of samples used for model updates. Initially, the unsupervised model $M$ with random initialized is trained on a small unlabeled first-round training dataset $X_{0}$ (Algorithm 1 line 3). Then, use $M$ to compute and collect the loss of $X_{0}$, and compute the threshold $T_{1}$ (Algorithm 1 lines 4-7). Subsequently, the online training process comprises two stages: The initial phase: when the abnormal loss is insufficient, only $T_{1}$ is used to generate pseudo-labels (Algorithm 1 lines 11-20). The second phase: both $T_{1}$ and $T_{2}$ are used to generate pseudo-labels (Algorithm 1 lines 21-33). At regular intervals, $T_{1}$ and $T_{2}$ are recalculated and the model $M$ and $M^{\prime}$ is trained (Algorithm 1 lines 34-39). The Adaptive Threshold Computation algorithm (Algorithm 2) computes an adaptive threshold for sample losses. It fits the best distribution to the data (Algorithm 2 line 2) and computes the percentile of this distribution as the threshold (Algorithm 2 line 3).
III-C Adaptive Threshold Calculation Technique
When new network samples arrive, the behavioral patterns of network traffic may change. Consequently, the predefined thresholds for classifying network samples (e.g., normal or abnormal) need to be periodically updated. Therefore, a novel dynamic threshold calculation strategy that automatically learns from the ongoing behaviors and adapts the corresponding thresholds to newer coming data. The initial thresholds $T_{1}$ and $T_{2}$ are calculated based on a limited dataset (e.g., the data from the first day or randomly selected 1% from the dataset). We use these “normal” samples to train the unsupervised learning model. Initially, because we assume all samples are normal, we use only one threshold $T_{1}$. Then, in the second round of training, we use the trained unsupervised model to predict the loss values for new samples. These loss values are divided into normal and abnormal samples using the threshold $T_{1}$. When a sufficient number of abnormal samples are collected (e.g., when the number of abnormal samples reaches 500), we calculate a second threshold $T_{2}$. From the third round of training onward, we follow the process described in Section II for training and updating. In this way, the threshold for the unsupervised model $M$ can be automatically determined for ongoing network anomaly detection.
We use the network loss databases to determine the thresholds. Two thresholds, $T_{1}$ and $T_{2}$, are utilized to identify true positive and true negative network traffic, which can then be used to train the interpretable supervised learning model $M^{\prime}$ during the online phase. The thresholds are selected as follows:
  1. 1.
    Phase 1: The Probability-Probability (P-P) plot is used to find the most suitable loss distributions that fit the datasets. We utilize the CIC-Darknet2020 as an example to prove the effectiveness of the proposed threshold calculation technique. To show its generality, we also illustrate details of the threshold calculation on the CIC-DoHBrw-2020 dataset in Appendix A. Fig. 4 illustrates the empirical distribution of the losses obtained from the Adaptive NAD model against the best-fitting theoretical distributions. By comparing several classical loss distributions, the log normal distribution is selected as the optimal distribution to determine the thresholds.
  2. 2.
    Phase 2: MLE is used to predict the parameters of the log normal distribution by maximizing a likelihood function. The estimation of parameters $\mu$ and $\sigma$ using the maximum likelihood technique from a sample of $n$ observations, $X=(x_{1},x_{2},...,x_{n})$ being X the loss observations, is calculated as follows:
    Theorem 1.
    Let $X\sim\text{lognormal}(\mu,\sigma^{2})$. The maximum likelihood estimators under $n$ observations of the parameters are:
    • (i)
      $\hat{\mu}=\frac{\sum_{j}\ln(x_{j})}{n}$
    • (ii)
      $\hat{\sigma}^{2}=\frac{\sum_{j}(\ln(x_{j})-\hat{\mu})}{n}$
    Proof.
    $\begin{split}L(\mu,\sigma^{2}|X)&=\prod_{j=1}^{n}[f(x_{j}|\mu,\sigma^{2})],\\ &={(2\pi\sigma^{2})}^{-n/2}\prod_{j}{x_{j}}^{-1}\\ &exp(\sum_{j}-\frac{(log(x_{j})-\mu)^{2}}{2\sigma^{2}}).\end{split}$
    Applying it to Eq. (2) to have the log likelihood function $l$, which is calculated as:
    $l(\mu,\sigma^{2}|X)$
    $=-\frac{n}{2}\ln(2\pi\sigma^{2})-\sum_{j}\ln(x_{j})-\frac{\sum_{j}\ln(x_{j})^{2}}{2\sigma^{2}}$
    $\ \ \ \ +\frac{\sum_{j}\ln(x_{j})\mu}{\sigma^{2}}-\frac{n\mu^{2}}{2\sigma^{2}}.$
    Next, we maximize $l$ by applying the derivative in terms of $\mu$ and $\sigma$. $\hat{\mu}$ value is obtained by:
    $\frac{\partial l}{\partial\mu}=\frac{\sum_{j}\ln(x_{j})}{\sigma^{2}}-\frac{n\hat{\mu}}{\sigma^{2}}=0,$
    $\hat{\mu}=\frac{\sum_{j}\ln(x_{j})}{n}.$
    The value of $\hat{\sigma}$ is calculated using the following:
    $\frac{\partial l}{\partial\sigma^{2}}=-\frac{\sum_{j}(\ln(x_{j})-\hat{\mu})^{2}}{2(\hat{\sigma}^{2})^{2}}+\frac{\sum_{j}(\ln(x_{j})-\hat{\mu})^{2}}{2(\hat{\sigma}^{2})^{2}}.$
    Then, letting $\frac{\partial l}{\partial\sigma^{2}}=0$, we have $\hat{\sigma}^{2}$ as:
    $\hat{\sigma}^{2}=\frac{\sum_{j}(\ln(x_{j})-\hat{\mu})}{n}.$
  3. 3.
    Phase 3:
    (a) Normal Distribution.
    (b) Abnormal Distribution.
    Figure 4: Loss distributions of normal and abnormal network traffic for CIC-Darknet2020.
    (a) Network Distributions.
    (b) Threshold update.
    Figure 5: Adaptive threshold calculation on CIC-Darknet2020.
    The threshold $T_{i}$ is obtained by Proposition 1.
    Proposition 1.
    Let $T_{1}$ and $T_{2}$ be the threshold for network traffic, and $p_{1}$ and $p_{2}$ are the percentiles for normal and abnormal loss distributions related to $T_{1}$ and $T_{2}$. If the normal and abnormal loss follows a log normal distribution with parameters $\mu_{i}$ and $\sigma_{i}$, being $i=1$ for normal and $i=2$ for abnormal traffic, then $T_{1}$ and $T_{2}$ are calculated as follows with a ($1-\alpha$) confidence level ($10<\alpha<$ 1):
    $T_{1}=exp(\Phi^{-1}(p_{1})\sigma_{1}+\mu_{1}),$
    $T_{2}=exp(\Phi^{-1}(1-p_{2})\sigma_{2}+\mu_{2}).$
    Proof.
    Let $X_{1}$ and $X_{2}$ be the loss of normal observations and loss of abnormal observation with cumulative distribution functions $F_{1}$ and $F_{2}$ respectively formulated as follows:
    $F_{1}(T_{1})=P(X_{1}\leq T_{1})=p_{1},$
    $F_{2}(T_{2})=P(X_{2}\geq T_{2})=p_{2}.$
    By assumption, $F_{i}(T_{i})$ follows a log normal distribution that is defined as follows:
    $F_{1}(T_{1})$ $=P(X_{1}\leq T_{1})$
[…]
    $F_{2}(T_{2})$ $=1-P(X_{2}\leq T_{2})$
[…]
    where $\Phi$ is the cumulative distribution function of the standard normal distribution, i.e., $N(0,1)$. $f_{1}(x_{1})$ and $f_{2}(x_{2})$ are the probability density function of the log normal distribution, which is calculated by:
    $f_{1}(x_{1})=f_{2}(x_{2})=\frac{1}{x\sqrt{2\pi}\sigma}exp(-\frac{(log(x)-\mu)^{2}}{2\sigma^{2}}).$
    Based on the above equations, the threshold $T_{1}$ and $T_{2}$ are calculated as follows:
    $p_{1}=\Phi\left(\frac{log(T_{1})-\mu_{1}}{\sigma_{1}}\right),$
    $p_{2}=1-\Phi\left(\frac{log(T_{2})-\mu_{2}}{\sigma_{2}}\right),$
    $\Phi^{-1}(p_{1})=\frac{log(T_{1})-\mu_{1}}{\sigma_{1}},$
    $\Phi^{-1}(1-p_{2})=\frac{log(T_{2})-\mu_{2}}{\sigma_{2}},$
    $T_{1}=exp(\Phi^{-1}(p_{1})\sigma_{1}+\mu_{1}),$
    $T_{2}=exp(\Phi^{-1}(1-p_{2})\sigma_{2}+\mu_{2}).$
In our model, the thresholds $T_{1}$ and $T_{2}$ are determined based on the quantiles of the best-fit statistical distributions for normal and abnormal losses, respectively. Specifically, $T_{1}$ should be set at a high percentile of the best-fit distribution for normal losses, while $T_{2}$ should be set at a low percentile of the best-fit distribution for abnormal losses. Selecting a high percentile for $T_{1}$ ensures that most normal samples, which typically have lower loss values, fall below it, thereby reducing false positives. Conversely, setting $T_{2}$ at the low percentile of the best-fit distribution for abnormal losses ensures that most abnormal samples, which typically have higher loss values, exceed it, minimizing false negatives.
The description of renewed thresholds of the proposed adaptive NAD model for both $T_{1}$ and $T_{2}$ over 8 renewal times is given in Fig. 5(b), where up to 5000 normal and 5000 abnormal loss values are retained in every renewal time and the old ones are discarded as new ones are added, ensuring that the most recent samples are always maintained. This approach allows the model to use the most up-to-date data for accurately determining the thresholds $T_{1}$ and $T_{2}$. We observe that the recent characteristics of network traffic determine threshold variations and $M$’s loss value is from 50 to 230. In addition, the thresholds for $T_{1}$ and $T_{2}$ ranges from 120 - 122 and 126 - 128, respectively.
III-D Complexity and storage analysis
Model complexity. The overall model complexity of Adaptive NAD is determined by the most complex component of the model. For instance, in this work, Adaptive NAD uses the threshold computation technique and two models, including LSTM-VAE (with an input sequence length of 5, hidden size of 64, latent size of 32, and a single layer) and RF. The LSTM-VAE is the most complex model, with 0.26M FLOPs and 58,207 parameters, which defines the model complexity of Adaptive NAD.
Storage cost. As for the space complexity, the primary cost includes two loss databases with fixed-size (e.g., 5000 in this work), FIFO queues, a temporary database from each batch that is periodically cleared, and all anomalies to assist security analysis for network operators.
IV Evaluation
This section presents the experimental setup and results. The datasets are detailed subsection IV-A. Next, the experiment setup is introduced in subsection IV-B. Finally, in subsection IV-C, we compare the eleven benchmarks and analyze their results.
IV-A Dataset
The CIC-Darknet2020 [27] network security dataset is employed to evaluate Adaptive NAD, which is a two-layered dataset (ISCXTor2017 and ISCXVPN2016) including both VPN and Tor encrypted traffic provided by Darknet’s hidden services. This paper evaluates Adaptive NAD on the Tor and non-Tor binary classification task, which has 93,309 normal and 1,392 abnormal entities over 13 days. We split the dataset into three subsets, i.e., a first-round training set (610 normal and 34 abnormal data) to train the initial unsupervised deep learning model of the first layer in the adaptive NAD framework so that it can generate high-confidence pseudo labels for the next rounds, a training set (85,165 normal and 1292 abnormal data) for online learning to update Adaptive NAD automatically, and testing set (7534 normal and 66 abnormal data) to evaluate and compare with state-of-the-art, where Day 2’s data is utilized as the first-round training set as Day 1’s data doesn’t have any normal data, day 1, as well as day 3-12’s data, is used as the training set, and day 13’s data is employed as the testing set.
In parallel, the CIC-DoHBrw-2020 [28] dataset is utilized to further evaluate the proposed Adaptive NAD’s performance. This dataset captures both benign and malicious DNS over HTTPS (DoH) traffic, as well as non-DoH traffic. This paper evaluates Adaptive NAD on the DoH and non-DoH binary classification task, which has 897,621 normal and 269,640 abnormal entities. The dataset was divided into three subsets: a first-round training set (1% of the total data, with 8,985 normal and 2,683 abnormal data), a training set (69% of the total data, with 622,045 normal and 186,869 abnormal data), and a testing set (30% of the total data, with 266,591 normal and 80,088 abnormal data).
The Edge-IIoTset dataset [29] is designed to evaluate intrusion detection systems (IDS) in IoT and IIoT environments. It is generated from a purpose-built testbed featuring diverse devices, sensors, protocols, and cloud/edge configurations. The dataset includes fourteen attack types related to IoT and IIoT connectivity protocols. This paper evaluates Adaptive NAD on the benign traffic and attack traffic binary classification task. which has 2,154,864 benign and 1,076,200 attack entities. The dataset was divided into three subsets: a first-round training set (1% of the total data, with 21,548 benign and 10,762 attack data), a training set (69% of the total data, with 1,486,856 benign and 742,578 attack data), and a testing set (30% of the total data, with 646,460 benign and 322,860 attack data).
The online training of Adaptive NAD does not use any label knowledge of legitimate and malicious data, which is a reasonable assumption because normal samples far outnumber anomalous ones in anomaly detection applications. This is why the first-round of the training set includes both normal samples and a small number of abnormal samples, without requiring any label information for either type of data.
IV-B Setup
System settings. Our NAD setup includes an encoder and a decoder, where both the encoder and the decoder have one LSTM layer with 64 hidden neurons, and the decoder has an additional fully connected (FC) layer. We use the Adam optimizer with a learning rate of 0.001. In the CIC-Darknet2020 [27], our timestep is 30 and update the model every 6,400 samples, while in both the CIC-DoHBrw-2020 [28] and Edge-IIoTset [29], our timestep is 5 and update the model in every 64,000 samples, which are the optimal configuration determined through extensive experiments. The input to NAD is network traffic data, and the latent dimension is 32. Based on our preliminary experiments, we calculated $T_{1}$ of the CIC-Darknet2020 [27] dataset using the 98th percentile and $T_{2}$ using the 10th percentile. For the CIC-DoHBrw-2020 [28] and Edge-IIoTset [29] datasets, we selected the 95th percentile to calculate $T_{1}$ and the 1st percentile to calculate $T_{2}$.
All simulations are carried out on a Linux server, with AMD EPYC 7282 16-Core Processor at 2.80 GHz and 377 GB of RAM. We implement the ADS using the Adaptive NAD framework in Python and construct the models with Pytorch and Scikit Learn libraries. All algorithms are implemented in the same environment, and the average results of 10 runs are collected for fair comparison.
Benchmarks. We use eleven benchmarks for comparison in this work to prove the performance of Adaptive NAD. Firstly, popular classical unsupervised machine learning algorithms, i.e., One-Class SVM (OCSVM) [33], Self-Organizing Maps (SOM) [33], Angle-Based Outlier Detector (AOD) [34], Isolation Forest (IF) [35], and KMeans [36], are used as offline candidates to compare the performance of Adaptive NAD. A Gaussian distribution for the latent variable $z$ is assumed.
Then, Adaptive NAD is compared with the state-of-the-art including GDN [14], where the relationship between different sensors is considered for improving detection performance and providing interpretability, DAGMM [22] achieving SPAUC improvement by jointly minimizing a combination of reconstruction loss and energy function, AOC-IDS [26] improves pseudo label generation with associate probabilities derived from normal and abnormal Gaussian distributions, DeepFT [11] with a dynamic threshold that is updated based on the extreme values observed in the input time-series data, DeepAID [15] employing a multivariate LSTM for anomaly detection by minimizing prediction loss, and OOF [21] which updated the threshold in a drifting manner and used the Mann-Whitney U test.
Afterward, we compared the performance of three different training settings: online training, which continuously updates the supervised and unsupervised models and thresholds with new data; initial online training, which trains the unsupervised model only with the first-round of the training dataset; and offline training, utilizing the entire training dataset to train the unsupervised model without further updates.
TABLE I: NAD, Fixed-NAD, and Adaptive NAD comparison on CIC-Darknet2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
NAD 84.88 52.68 91.62 50.97 15.24 1.52 57.00
Fixed-NAD 90.60 54.00 92.26 54.93 9.43 6.06 61.48
Adaptive NAD 99.89 96.13 98.07 96.90 0.08 3.80 97.70
TABLE II: NAD, Fixed-NAD, and Adaptive NAD comparison on CIC-DoHBrw-2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
NAD 83.24 89.45 64.07 67.05 0.30 71.55 63.69
Fixed-NAD 84.03 90.67 65.61 69.08 0.15 68.63 65.40
Adaptive NAD 93.22 92.63 87.79 89.90 2.12 22.30 80.21
TABLE III: NAD, Fixed-NAD, and Adaptive NAD comparison on Edge-IIoTset.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
NAD 66.77 83.24 50.11 40.26 0.003 99.76 50.11
Fixed-NAD 79.28 77.40 74.29 75.39 10.77 40.63 55.78
Adaptive NAD 99.45 99.58 99.20 99.38 0.03 1.57 99.06
TABLE IV: Comparison with classical unsupervised machine learning algorithms on CIC-Darknet2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
OCSVM 68.92 49.44 36.26 40.88 30.50 96.97 48.8
SOM 89.10 49.03 46.09 47.35 9.07 98.75 48.90
AOD 71.16 49.59 40.39 41.84 90.91 0.66 49.13
IF 57.58 50.90 75.60 38.51 42.74 6.06 57.01
KMeans 60.90 49.43 34.47 38.01 38.64 92.42 48.97
Adaptive NAD 99.89 96.13 98.07 96.90 0.08 3.80 97.70
TABLE V: Comparison with classical unsupervised machine learning algorithms on CIC-DoHBrw-2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
OCSVM 51.50 51.50 52.11 47.71 49.03 46.76 50.11
SOM 77.54 66.77 59.83 61.01 7.26 73.08 53.47
AOD 69.68 44.48 47.25 44.84 11.07 94.42 49.36
IF 77.88 67.67 58.95 59.96 5.86 76.23 53.91
KMeans 89.97 85.81 85.82 86.04 6.66 21.27 63.88
Adaptive NAD 93.22 92.63 87.79 89.90 2.12 22.30 80.21
TABLE VI: Comparison with classical unsupervised machine learning algorithms on Edge-IIoTset.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
OCSVM 83.18 89.63 74.84 77.58 0.17 50.14 74.45
SOM 57.55 36.43 43.92 38.67 15.26 96.89 48.98
AOD 67.25 74.81 75.24 67.25 48.68 0.83 51.33
IF 68.65 83.94 52.93 46.02 0.004 94.13 52.93
KMeans 50.40 38.61 40.66 39.31 30.17 88.51 49.21
Adaptive NAD 99.46 99.58 99.20 99.38 0.03 1.57 99.06
Evaluation Metrics. We use Accuracy (Acc.), Precision-macro (Pre.), Recall-macro (Rec.), F1-score-macro (F1.), False Alarm Rate (FAR), Missed Detection Rate (MDR), and standardized partial AUC (SPAUC) [37] to measure different models’ performance. Particularly, SPAUC measures the performance of a model within a specific region of the ROC curve, focusing on a segment that is most relevant under conditions of class imbalance and differential costs of misclassification. FAR reflects the amount of normal network traffic that is wrongly classified as abnormal while MDR indicates how much anomaly network traffic that are miss-detected by the models. The higher the SPAUC, the better the performance is. The lower the FAR and MDR, the better the performance is.
Here we use $\textit{SPAUC}_{\textit{FPR}\leq 0.05}$ for all experiments. The number of layers of the comparative models like GDN, DAGMM, AOC-IDS, DeepFT, DeepAID, and OOF are 2, 4, 4, 6, 2, and 4, respectively. As offline classical machine learning approaches need to build models in the offline phase, both the first-round training and training sets are selected to train these models, and the testing set is utilized to evaluate the models.
IV-C Results and Analysis
IV-C1 Baseline Comparison Summary
This section introduces the comparison of RF with some classical machine learning methods firstly. Then, the contribution of each key technique in Adaptive NAD is verified.
Comparisons with RF. Fig. 6(a) illustrates that RF outperforms other classic machine learning models like SVM, KNN, and GNB by 15.04%, 3.76%, and 37.93%, demonstrating that RF is better for being $M^{\prime}$ on the second phase. Besides, from the right subfigure in Fig. 6(b) we can see that the optimal number of estimators of RF is 40 on the CIC-Darknet2020 dataset. Similar performance has been observed on the CIC-DoHBrw-2020 and Edge-IIoTset datasets, and the optimal number of estimators of RF is 40 also.
SVMKNNGNBRF0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111SPAUC
(a) Supervised model comparison.
10101010202020203030303040404040505050508080808090909090100100100100Number of EstimatorsSPAUC (%)
(b) Different RF parameter evaluation.
Figure 6: Performance comparison over different supervised machine learning models on CIC-Darknet2020 and the effect of a different number of estimators of RF on SPAUC.
Contribution of each key technique. Table I, Table II, and Table III illustrate the contribution of each key technique, where NAD means that only the unsupervised deep learning $M^{\prime}$ is utilized, and Fixed-NAD is a variant of Adaptive NAD with a constant threshold.
As shown in Table I, Table II, and Table III, Adaptive NAD outperforms other methods. Adaptive NAD achieves detection SPAUC as high as 97.70%, 80.21%, and 99.06% on CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets with about 36.22%, 14.81%, and 43.28% improvement compared to Fixed-NAD and 40.70%, 13.48%, and 45.95% improvement compared with NAD, respectively. In addition, The MDR of NAD is 2.28% lower but its FAR is 15.16% higher than that of Adaptive NAD on the CIC-Darknet2020 dataset, indicating that it balances FAR and MDR less effectively than Adaptive NAD. Then, NAD and Fixed-NAD exhibited a lower FAR on CIC-DoHBrw-2020. However, their MDRs are very high, at 71.55% and 68.63%, respectively. This indicates that they will miss a large number of anomalies. Notably, both the FAR and MDR of the proposed Adaptive NAD are far lower than NAD and Fixed-NAD on the Edge-IIoTset dataset, proving the effectiveness of Adaptive NAD.
IV-C2 Comparison with state-of-the-art
This section compares the performance of Adaptive NAD with the state-of-the-art. We rely primarily on FAR, MDR, and SPAUC as they can better represent anomalous behavior in ADSs.
Comparison with unsupervised machine learning models. As shown in Table IV, Table V, and Table VI, Adaptive NAD outperforms classical unsupervised machine learning algorithms on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets. It achieves the highest detection SPAUC of 97.70% on the CIC-Darknet2020 dataset with an improvement of approximately 40.69%, 80.21% on the CIC-DoHBrw-2020 dataset with an improvement of approximately 16.33%, and 99.06% on the Edge-IIoTset dataset with an improvement of approximately 24.61%. Adaptive NAD exhibits excellent FAR while maintaining a very low MDR on all the datasets. Specifically, it achieves the lowest FAR of 0.08%, 2.12%, and 0.03% on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively, meaning it generates very few false alarms while accurately detecting anomalies.
Comparison with recent advanced unsupervised algorithms. Table VII, Table VIII, and Table IX shows Adaptive NAD compared with the state-of-the-art unsupervised models in respect of different metrics for the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively. Table VII demonstrates Adaptive NAD surpasses all comparison methods across all metrics on the CIC-Darknet2020 dataset with a remarkable SPAUC of over 97.70% and the lowest FAR of 0.08%. This high level of performance highlights Adaptive NAD’s capacity in anomaly detection within dynamic environments. In Table VIII, Adaptive NAD achieves the highest detection SPAUC of 80.21%, improving by approximately 23.04%. It also performs the best in terms of FAR and MDR, exhibiting the lowest MDR and the third-lowest FAR. Although DeepAID and DAGMM have lower FARs, their MDRs are significantly higher, with DeepAID’s MDR as high as 99.76% and DAGMM’s MDR at 83.82%, meaning that these models fail to detect most actual anomalies. Table IX depicts that Adaptive NAD obtains the highest Accuracy, Precision, Recall, F1-score, MDR, and SPAUC on the Edge-IIoTset dataset, indicating its superior anomaly detection capabilities in comprehensive realistic IoT and IIoT cyber security scenarios. Although its FAR is higher than DeepAID, it has a 9.92% lower MDR and 49.65% higher SPAUC compared with DeepAID. Overall, considering all the metrics including accuracy, precision-macro, recall-macro, f1-score-macro, FAR, MDR, and SPAUC, Adaptive NAD is the best design for network anomaly detection.
TABLE VII: Comparison with the state-of-the-art unsupervised models on CIC-Darknet2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
GDN 40.77 49.93 49.89 38.51 63.90 36.32 50.00
DAGMM 95.86 64.31 83.65 69.64 3.61 29.09 71.98
AOC-IDS 41.97 50.60 67.00 30.68 58.47 7.58 50.74
DeepFT 69.94 60.66 62.85 61.20 24.26 50.04 51.36
DeepAID 99.48 93.67 93.39 93.53 0.26 12.96 92.35
OOF 94.72 57.42 95.34 64.42 5.30 4.04 73.81
Adaptive NAD 99.89 96.13 98.07 96.90 0.08 3.80 97.70
TABLE VIII: Comparison with the state-of-the-art unsupervised models on CIC-DoHBrw-2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
GDN 86.32 76.50 82.50 78.82 11.62 23.37 57.17
DAGMM 79.38 77.19 57.27 57.30 1.64 83.82 56.15
AOC-IDS 61.26 53.37 54.30 52.89 32.76 58.63 50.34
DeepFT 49.04 47.64 46.62 43.98 48.99 57.76 49.82
DeepAID 76.58 55.10 50.04 43.69 0.14 99.76 50.05
OOF 78.92 72.37 77.85 73.89 20.17 24.12 53.54
Adaptive NAD 93.22 92.63 87.79 89.90 2.12 22.30 80.21
TABLE IX: Comparison with the state-of-the-art unsupervised models on Edge-IIoTset.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
GDN 99.42 99.42 99.28 99.35 0.58 0.56 96.77
DAGMM 90.34 51.35 95.15 50.10 4.84 4.84 61.95
AOC-IDS 84.25 86.40 85.54 84.23 1.69 27.21 79.89
DeepFT 45.03 43.33 46.51 38.56 87.93 19.03 49.89
DeepAID 66.89 82.91 50.29 40.64 0.003 9.94 50.29
OOF 74.16 77.31 80.04 73.94 37.55 2.36 52.05
Adaptive NAD 99.99 99.99 99.99 99.99 0.39 0.016 99.97
IV-C3 Comparison with different training settings (e.g., online, initial, and offline)
This section compares the performance of three different training settings—online training, initial online training, and offline training—regarding the average metrics over two varied public datasets. We rely primarily on FAR, MDR, and SPAUC as they can better represent anomaly detection performance in ADSs.
As shown in Table X, Table XI, and Table XII, online training outperforms other training settings on all three datasets, it achieves SPAUC as high as 97.70%, 80.21%, and 99.06% on CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets with about 11.37%, 16.7%, 47.67% improvement, respectively. In the CIC-Darknet2020 dataset, offline training achieved the lowest MDR. However, it has a relatively high FAR of 27.74%, which means it is prone to generating a significant number of false alarms. Then, in the CIC-DoHBrw-2020 dataset, initial online training and offline training achieved the lowest MDR and FAR, respectively. However, the FAR of initial online training is 7.10%, nearly five percent higher than online training, while the MDR of offline training is as high as 80.10%. Overall, online training balances FAR and MDR better. In the Edge-IIoTset dataset, online training performs best across all evaluation metrics. In particular, it has 47.67% and 26.35% higher SPAUC than initial online training and offline training, proving the proposed Adaptive NAD’s superior performance for network anomaly detection.
Intuitively, the performance of initial online training is better than that of offline training because initial online training uses the first-round training dataset, which does not contain enough anomalous samples to impact the unsupervised model. In contrast, offline training uses the entire training dataset, which includes a large number of anomalous samples, adversely affecting the model’s performance.
TABLE X: Performance (%) of Adaptive NAD on CIC-Darkent2020: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
Setting Acc. Pre. Rec. F1. FAR MDR SPAUC
Online 99.89 96.13 98.07 96.90 0.08 3.80 97.70
Initial 99.58 88.17 87.02 87.59 0.20 25.76 86.33
$\triangledown$ +0.31 +7.96 +11.05 +9.31 -0.12 -21.96 +11.37
Offline 72.49 51.50 85.37 44.88 27.74 1.52 53.27
$\triangledown$ +27.4 +44.63 +12.7 +52.02 -27.66 +2.28 +44.43
TABLE XI: Performance (%) of Adaptive NAD on CIC-DoHBrw-2020: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
Setting Acc. Pre. Rec. F1. FAR MDR SPAUC
Online 93.22 92.63 87.79 89.90 2.12 22.30 80.21
Initial 90.33 86.01 87.33 86.64 7.10 18.27 63.51
$\triangledown$ +2.89 +6.62 +0.46 +3.26 +4.98 -4.03 +16.70
Offline 81.00 85.30 59.63 60.78 0.66 80.10 59.04
$\triangledown$ +12.22 +7.33 +28.16 +29.12 +1.46 -57.80 +21.17
TABLE XII: Performance (%) of Adaptive NAD on Edge-IIoTset: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
Setting Acc. Pre. Rec. F1. FAR MDR SPAUC
Online 99.45 99.58 99.20 99.38 0.03 1.57 99.06
Initial 67.51 73.03 51.45 43.39 0.45 96.64 51.39
$\triangledown$ +31.94 +26.55 +47.75 +55.99 -0.42 -95.07 +47.67
Offline 81.91 89.18 72.87 75.41 0.07 54.18 72.71
$\triangledown$ +17.54 +10.04 +26.33 +23.91 -0.04 -52.61 +26.35
V Discussions
This section discusses the main design choices, limitations, and future works.
Why Adaptive NAD is unsupervised? The online training of Adaptive NAD does not utilize any label knowledge of normal events or anomalies, making it a fully unsupervised framework. In the first-round training phase, a limited subset of network events (approximately 1% of the total data) without label information is used to train the unsupervised deep learning model, which constitutes the first layer of the Adaptive NAD framework (see Section II). Following this, Adaptive NAD is deployed for online learning with incoming data, still without using any label information.
In this work, we assume that anomalies are significantly fewer in number compared to normal events, which is a reasonable assumption in many practical security-related applications. Based on this assumption, the trained model in the first layer of the Adaptive NAD framework can generate a few high-confidence pseudo labels without any ground truth labels.
Our experimental results also demonstrate that, with a small amount of anomalous noise in the first-round training set, the unsupervised deep learning model of the first layer can still learn the patterns of normal samples and provide a few reliable pseudo-labels. However, when the proportion of anomalous samples in the first-round training set is high, the unsupervised model tends to fail to train effectively and cannot produce enough reliable pseudo-labels. In such cases, a certain number of normal samples with ground truth labels are expected to train the unsupervised model in the Adaptive NAD’s first layer.
Why Adaptive NAD is interpretable? The interpretable supervised machine learning models (e.g., Random Forest, Decision Trees, and Linear models) can be chosen as the second layer of the proposed Adaptive NAD to output the final prediction results. Thus, all results predicted by Adaptive NAD are interpretable.
As we all know, deep learning models can extract high-quality patterns and accomplish good results in complex network environments. However, in practical network security applications, deep learning models (e.g., LSTM) for anomaly detection are not easy to deploy in a real-world system. Since the models are not transparent and their prediction results are not interpretable, it is hard to trust the system’s decision (e.g., normal events or anomalies) without sufficient reasons and credible evidence. In this work, the proposed Adaptive NAD framework is a fully unsupervised interpretable model to assist operators in understanding the models’ decisions and increase trustworthiness for the prediction results, which is critical for improving the practicality of security systems in the real world.
Limitations and future work. Firstly, LSTM-VAE and RF are used in the current framework. More models could be attempted in the future to select as first or second layer of Adaptive NAD for optimizing the performance of the proposed framework. Secondly, we configure the hyper-parameters of Adaptive NAD empirically, which lacks systematic strategies to configure them. Future work can investigate more techniques to configure the hyper-parameters. Thirdly, three public datasets are employed to verify the proposed model. In the future, we intend to gather and integrate network events from different real-world systems and use these data to further evaluate the robustness of ADSs built using the Adaptive NAD framework. Fourthly, in addition to the applications in the security domain, our work could be generalized to other domains for detecting outliers.
VI Related Work
In this section, we briefly discuss existing literature from the following three aspects, unsupervised anomaly detection models and online unsupervised anomaly detection approaches.
Unsupervised anomaly detection models: The ”Zero-positive” setting, where some normal samples are known by a model, is one of the most common assumptions for unsupervised anomaly detection. Existing literature from this setting can be roughly grouped into three categories, i.e., reconstruction-based, prediction-based, and contrastive-based approaches.
Reconstruction-based methods use the encoder and decoder to reconstruct the input and those with large reconstruction errors are recognized as anomalies. For instance, [8] trains their models by minimizing the distance between original normal data and reconstructed normal data, and anomalies are recognized when the reconstruction error of a sample is high. [9] leverages a Variational Autoencoder (VAE) to model normal seasonal KPI patterns without labels, enabling anomaly detection through reconstruction probability analysis in latent space. Additionally, instead of using one encoder and one decoder as in [8, 9], [10, 11] utilizes multiple encoders and decoders to enhance the representation ability of the model. Moreover, in [13], spectral analysis is employed within a pre-trained autoencoder to synchronize phase information across complex asynchronous multivariate time series data for anomaly detection by minimizing quantile reconstruction losses. [12] introduces an unsupervised approach that models normal data through a hierarchical Variational AutoEncoder with two latent variables for capturing inter-metric and temporal dependencies in time series data. Reconstruction-based methods do not demand any labeled anomalies and are well-suited for non-time series data. For time-series data, prediction-based and contrastive-based methods are usually utilized, whereas prediction-based approaches usually use the past series to forecast future windows or specific time stamps while contrastive-based methods learn data representations through comparisons between positive and negative samples. Given an intuitive assumption, which is neighboring time windows or timestamps in the raw time series share high similarity, thus are considered as positive samples, and those from distant time windows or timestamps are treated as negative samples [38]. Similar to reconstruction-based methods, prediction-based approaches can predict those from normal time series with higher probability while tending to wrongly predict abnormal data. [18] considers the scenarios in which operators can manually examine a few suspicious data samples to provide their labels. It improves anomaly detection performance by maximizing the prediction loss to unlearn the reported abnormal samples. [14] improves anomaly detection performance by considering relationships between different sensors. It also provides interpretability through the learned graph edges and attention weights, indicating the importance of relationships between sensors. [11] generates self-supervised anomaly labels by reconstructing an input time-series window using a system state decoder, which provides an accurate prediction of the subsequent state. Anomalous labels are assigned to instances where the predicted state exceeds a dynamically determined threshold, while benign labels are assigned to instances where the predicted state falls below this threshold. The threshold is dynamically adjusted based on the extreme values observed in the input time-series data. [15] employs a typical prediction-based anomaly detection method that trains a multivariate LSTM by minimizing prediction loss of the model’s performance of predicting the current time series using previous time series data. As a contrastive-based method, [39] proposes a contrastive learning-based method that learns representations that distinguish normal time series points from anomalies by maximizing the similarity between two views (patch-wise and in-patch representations) of normal data and enlarging the discrepancy for anomalies. [40] proposed a self-supervised framework tailored for non-stationary time series based on positive-unlabeled (PU) learning that treats negative samples as unknown samples and assigns weights to these samples to handle the problem of sampling bias incurred by non-stationary characteristics of most time series.
Except for these, some research uses hybrid approaches that optimize a combination of reconstruction loss, prediction loss, or contrastive loss. For instance, [22] combines a deep autoencoder with a Gaussian Mixture Model (GMM) and detects anomalies by jointly minimizing a combination of reconstruction loss and energy function. It outperforms traditional clustering methods (e.g., k-means) by effectively handling high-dimensional data through integrated dimensionality reduction and clustering in a single step. [16] develops a Fused Sparse Autoencoder and Graph Net, which minimizes reconstruction and prediction losses while modeling the relationships between features in multivariate time series. [17] introduces a self-supervised learning framework that learns time series representations by jointly optimizing reconstruction loss and contrastive loss for both segment and instance levels. It also enhances contrastive learning’s performance by injecting more hard negative samples.
Although many previous works have validated the flexibility and effectiveness of ”zero-positive” anomaly detection through reconstruction-based, prediction-based, or contrastive learning-based approaches, essential aspects that are still not fully studied are the training of the initial model without any assumptions of labeled data, including normal ones, and the effective incremental update of the initial model to changing new patterns without any label supervision. Besides, most existing work lacks interpretability or struggles with high false alarm rates or long latency, while those features are vital for many practical applications such as smart healthcare, manufacturing, or autonomous driving [1, 2, 3]. Hence, unlike the aforementioned literature, we design a practical self-adaptive unsupervised anomaly-detection framework without assuming any prior knowledge of the network traffic and can effectively adaptively and incrementally update the model.
Online unsupervised anomaly detection approaches: Recently, there is only limited research considering online unsupervised anomaly detection for improving ADSs’ practicality, such as reducing the false alarm rate [23, 24, 25, 8, 18, 26, 21, 20], detecting concept drift [8, 18, 26, 21, 20], and providing interpretability [25, 15]. However, most of them like [23, 24, 25] consider “online” as the model being trained offline and utilized for real-time inferences without model updating. Other research like [8, 18, 26, 21, 20] proposes real online unsupervised anomaly detection methods that incrementally update the model to remain synchronized with dynamic environments. For example, [26] introduces a pseudo-label generator that generates data labels in a streaming fashion to re-train the model. [21] adapts the threshold when a batch of input data varies from normal data. However, most existing studies assume known normal or abnormal samples and research on incremental model updates without any label supervision is still open. Besides, most existing work struggles with high FAR and lacks interpretability, while these metrics are vital for practical applications such as smart healthcare, manufacturing, or autonomous driving [1, 2, 3].
VII Conclusions
In this paper, we propose Adaptive NAD, a general framework to improve and interpret online unsupervised anomaly detection in real-world applications in security domains. In Adaptive NAD, an interpretable two-layer anomaly detection strategy (ITAD-S) is designed to generate reliable high-confidence pseudo-labels and provide generalization and interpretability. Besides, an online learning scheme has been developed to update both the unsupervised deep learning model and the loss distributions using a well-designed threshold calculation technique that offers low false positives and adapts to new threats. By applying and evaluating Adaptive NAD over two classic anomaly detection datasets, we demonstrate that Adaptive NAD can provide human-readable, low false positive rates and can adapt to new threats online anomaly detection results.
Appendix A
To show the generality of the proposed threshold calculation technique, we also evaluate it on the CIC-DoHBrw-2020 dataset. Fig. 7 is the empirical distribution of the losses obtained from the proposed Adaptive NAD against the best-fitting theoretical distributions. From Fig. 7, we can see that the norm and logistic distributions are the best-fitting theoretical distributions of the normal and abnormal losses.
Here, MLE and Method of Moments (MoM) are utilized to estimate the parameters of norm and logistic distributions respectively. The parameters of the norm distribution are calculated as follows:
Theorem 2.
Let $X\sim\text{normal}(\mu,\sigma^{2})$. The maximum likelihood estimators under $n$ observations of the parameters are:
  • (i)
    $\hat{\mu}=\frac{\sum_{j=1}^{n}x_{j}}{n},$
  • (ii)
    $\hat{\sigma^{2}}=\frac{\sum_{j=1}^{n}(x_{j}-\hat{\mu})}{n}$
Proof.
$\begin{split}L(\mu,\sigma^{2}|X)&=\prod_{j=1}^{n}[f(x_{j}|\mu,\sigma^{2})],\\ &=\prod_{j=1}^{n}\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x_{j}-\mu)^{2}}{2\sigma^{2}}},\\ &={(2\pi\sigma^{2})}^{-n/2}e^{-\frac{(x_{j}-\mu)^{2}}{2\sigma^{2}}},\\ &={-\frac{n}{2}ln(2\pi\sigma^{2})-\sum_{j=1}^{n}{\frac{(x_{j}-\mu)^{2}}{2\sigma^{2}}}}.\end{split}$
Let:
$\frac{\partial L}{\partial\mu}=\frac{1}{\sigma^{2}}\sum_{j=1}^{n}(x_{j}-\mu)=0,$
we have $\hat{\mu}$ as:
$\hat{\mu}=\frac{\sum_{j=1}^{n}x_{j}}{n}.$
Let
$\frac{\partial L}{\partial\sigma^{2}}=-\frac{n}{2}\cdot\frac{1}{\sigma^{2}}+\frac{1}{2\sigma^{4}}\sum_{j=1}^{n}(x_{j}-\mu)^{2},$
we have $\hat{\sigma^{2}}$ as:
$\hat{\sigma^{2}}=\frac{1}{n}\sum_{j=1}^{n}(x_{j}-\frac{\sum_{j=1}^{n}x_{j}}{n})^{2}.$
The parameter estimation of the logistic distribution is below:
Theorem 3.
Let $X\sim\text{logistic}(\mu,\gamma)$. The maximum likelihood estimators under $n$ observations of the parameters are:
  • (i)
    $\hat{\mu}=\frac{1}{n}\sum_{j=1}^{n}x_{j},$
  • (ii)
    $\hat{\gamma}=\frac{1}{\pi}\sqrt{\frac{3}{n}\cdot\sum_{j=1}^{n}(x_{j}-\frac{\sum_{j=1}^{n}x_{j}}{n})^{2}}.$
Proof.
Because that the expectation value of the logistic distribution is $\mu$ and its variance is $\frac{\pi^{2}}{3}\gamma^{2}$, then
$\mu=\overline{x},$
which means:
$\hat{\mu}=\frac{1}{n}\sum_{j=1}^{n}x_{j}.$
Similarly,
$\frac{\pi^{2}}{3}\gamma^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{j}-\mu)^{2},$
$\hat{\gamma}=\frac{1}{\pi}\sqrt{\frac{3}{n}\cdot\sum_{j=1}^{n}(x_{j}-\frac{\sum_{j=1}^{n}x_{j}}{n})^{2}}.$
The threshold $T_{1}$ and $T_{1}$ for the CIC-DoHBrw-2020 dataset is obtained by Proposition 2.
Proposition 2.
Let $T_{1}$ and $T_{2}$ be the threshold for network traffic, and $p_{1}$ and $p_{2}$ are the percentiles for normal and abnormal loss distributions related to $T_{1}$ and $T_{2}$. If the normal loss follows a normal distribution with parameter $\mu_{1}$ and $\sigma$ while the abnormal loss follows a logistic distribution with parameter $\mu_{2}$ and $\gamma$, then $T_{1}$ and $T_{2}$ are calculated as follows with a ($1-\alpha$) confidence level ($10<\alpha<$ 1):
$T_{1}=\Phi_{1}^{-1}(p_{1})\sigma+\mu_{1}.$
$T_{2}=\Phi_{2}^{-1}(1-p_{2})\gamma+\mu_{2}.$
Proof.
Let $X_{1}$ and $X_{2}$ be the loss of normal observations and loss of abnormal observation with cumulative distribution functions $F_{1}$ and $F_{2}$ respectively formulated as follows:
$F_{1}(T_{1})=P(X_{1}\leq T_{1})=p_{1},$
$F_{2}(T_{2})=P(X_{2}\geq T_{2})=p_{2}.$
By assumption, $F_{1}(T_{1})$ follows a normal distribution and $F_{2}(T_{2})$ follows a logistic distribution that are defined as follows:
$F_{1}(T_{1})$ $=P(X_{1}\leq T_{1})$
[…]
$F_{2}(T_{2})$ $=1-P(X_{2}\leq T_{2})$
[…]
where $\Phi_{1}$ and $\Phi_{2}$ are the cumulative distribution functions of the normal and logistic distributions. $f_{1}(x_{1})$ and $f_{2}(x_{2})$ are the probability density function of the normal and logistic distributions, which are calculated by:
$f_{1}(x_{1})=\frac{1}{\sqrt{2\pi}\sigma}exp(-\frac{(x-\mu)^{2}}{2\sigma^{2}}),$ (8)
$f_{2}(x_{2})=\frac{e^{-\frac{x-u}{\gamma}}}{\gamma(1+e^{-\frac{x-u}{\gamma}})^{2}}.$
Based on the above equations, the threshold $T_{1}$ and $T_{2}$ are calculated as follows:
$p_{1}=\Phi_{1}\left(\frac{T_{1}-\mu_{1}}{\sigma}\right),$
$p_{2}=1-\Phi_{2}\left(\frac{T_{2}-\mu_{2}}{\gamma}\right),$
$\Phi_{1}^{-1}(p_{1})=\frac{T_{1}-\mu_{1}}{\sigma},$
$\Phi_{2}^{-1}(1-p_{2})=\frac{T_{2}-\mu_{2}}{\gamma},$
$T_{1}=\Phi_{1}^{-1}(p_{1})\sigma+\mu_{1},$
$T_{2}=\Phi_{2}^{-1}(1-p_{2})\gamma+\mu_{2}.$
(a) Normal Distribution.
(b) Abnormal Distribution.
Figure 7: Loss distributions of normal and abnormal network traffic for CIC-DoHBrw-2020.
(a) Network Distributions.
(b) Threshold update.
Figure 8: Adaptive threshold calculation on CIC-DoHBrw-2020.
Fig. 7 illustrates the empirical distribution of the losses obtained from the Adaptive NAD model against the best-fitting theoretical distributions on the CIC-DoHBrw-2020 dataset. The description of renewed thresholds of the proposed adaptive NAD model for both $T_{1}$ and $T_{2}$ on the CIC-DoHBrw-2020 dataset is given in Fig. 8.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62406215) and in part by the supported by the ”Fundamental Research Funds for the Central Universities”.